OpenStudy (anonymous):

f(x)=x^2+x and g(x)=2/(x+3) find value from : (f-g) (2)

OpenStudy (angela210793):

f-g=x^2+x-2/(x+3)=[(x^2+x)*(x-3)-2]/(x-3)=(x^3-2x^2-3x-2)/(x-3) f-g(2)=(8-8-6-2)/(2-3)=-8/-1=8

OpenStudy (anonymous):

how about it... f(x)=x^2+x and g(x)=2/(x+3) find value from : (f/g) (1)